# arXiv:2602.24140v1[physics.ins-det] 27 Feb 2026

### v1

@@ VI Conclusion
This paper has presented a radiation-tolerant current digitizer ASIC designed for beam loss monitoring applications in the HL-LHC environment. The architecture is derived directly from stringent system-level requirements, including a wide input dynamic range (1 mA to 1 pA), short integration windows (down to 10 µs), deterministic overload recovery, and tolerance to total ionizing doses exceeding 100 Mrad.
A first-order delta-sigma topology was selected to ensure unconditional stability, predictable saturation behavior, and robustness under large transient inputs, as required in machine protection systems. The achievable dynamic range scales with the product of sampling frequency and acquisition time, enabling high resolution for quasi-static signals while preserving fast response capability. In 10 µs mode, the system achieves 11-bit resolution, while extended acquisition enables significantly higher effective resolution for low-level signals.

### v2

@@ VI Conclusion
This paper has presented a radiation-tolerant current digitizer ASIC designed for beam loss monitoring applications in the HL-LHC environment. The architecture is derived directly from stringent system-level requirements, including a wide input dynamic range [1 mA, 1 pA], short integration windows down to 10 µs, deterministic overload recovery, and tolerance to total ionizing doses exceeding 100 Mrad.
A first-order delta-sigma topology was selected to ensure robust bounded behavior within its operational range, predictable saturation behavior, and robustness under large transient inputs, as required in machine protection systems. The achievable operating dynamic range, which is a function of the input current, scales with the acquisition time. This property allows the converter to deliver high resolution for quasi-static signals while maintaining the fast response required for transient events. In 10 µs mode, the system achieves 11 bit resolution, while extended acquisition enables significantly higher effective resolution.